OpenStudy (anonymous):

C=5/9(F-32) Solve for f. Its a literal equation

OpenStudy (anonymous):

c = 5/9 (f - 32) 5/9 (f - 32) = c Divide each side by 5/9. (Which is the same as multiplying by 9/5.) f - 32 = (9/5)c Add 32 to each side. f = (9/5)c + 32